OpenStudy (anonymous):

Michael bought 3/4 pound of sliced turkey. He ate 1/3 of the turkey. How much turkey did Michael eat?

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

it just \[\frac{ 3 }{ 4 } \times \frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

so... 1/4?

OpenStudy (anonymous):

yup